# Resolvi: A Reference Architecture for Scalable and Interoperable Entity Resolution

### v1

@@ 5.6 Clustering Engine
Clustering plays a crucial role in entity resolution (ER) by grouping linked entity references that correspond to the same real-world object. It follows the principle that things which co-occur or exhibit strong relational ties likely refer to the same underlying entity. This mechanism is complementary to matching because similarity as a general concept may provide the evidence of solid relational ties. Both mechanisms contribute to ER in an essential way, though only at least one of them is strictly necessary.
@@
Widely regarded as an essential step, not all reference systems implement it transparently [21, 33, 44]. Many entity matching systems and some batch ER systems, which operate on one or two data sources, focus solely on identifying matching record pairs, partially or completely omitting transitive resolution. In some cases, this approach can be beneficial, as it offloads the bad-triplet problem [1] to downstream systems.
@@
At the other end of the spectrum, entity alignment and record linkage systems designed around graph models inherently support clustering. Some ER solutions go further by eliminating the matching phase entirely. For example, clustering may subsume matching by employing functional similarity measures and a carefully chosen merge function[3], generic distance metrics[69], or probabilistic models that form clusters directly from the comparison space using similarity functions to model colocation [38].
@@
When applied, clustering typically operates on data supplied by the reference store. In cases where matching is omitted, the entity reference store effectively defines the relevant comparison space. Systems such as the JedAI Toolkit[51] and FAMER[58] already support various clustering algorithms [48, 59], offering flexible options for different ER workflows.
@@
Given its varying adoption across reference systems and the limited research on cluster coherence [68], system architectures may design a clustering engine if their specific problem domain and ER model of choice require one. The Strategy design pattern [19] may be used to encapsulate the various algorithms sitting at the core of the clustering engine.
New system architectures should carefully consider clustering engines that work with incremental ER. Only some clustering algorithms, such as agglomerative clustering, are generally incremental [24]. When a clustering engine is designed as part of a solution architecture, the architecture should incorporate designs for evaluating clustering performance, too. Few systems provide clustering evaluation, with d-blink [38] being one of the exceptions. Future efforts in this area would improve ER system transparency and reliability.
@@ 5.7 Presentation Layer
The logical component found across all ER systems is the entity profile presentation layer which has the responsibility of making entity profiles available to ER consumers. While entity profile formatting or display vary according to the needs of each implementation, all ER systems define some component which assembles entity profiles. The data structure used to represent entity profiles varies significantly across ER models, as different approaches to ER use distinct data representations. The act of designing ER system architectures necessarily involves choosing an ER model which best suits the problem domain. The entity profile assembler logic will necessarily be dictated by the selected ER model.
@@
In existing systems, common ways to represent entity profiles include pairs of matching entity references [33, 21], entity reference clusters [51, 38] or graphs [58]. Less commonly, in entity alignment [63], entity profiles may be defined based on relationships between attributes rather than between entity references.
SERF [3] introduces a black-box merge function that constructs more dominant entity references incrementally using the matching engine’s output. The R-Swoosh algorithm which defines this approach is implemented by some of the existing ER systems [65]. The final entity references constructed in this way represent the entity profiles, their data structure being determined flexibly by the black box merge function. Though not explicitly observed in practice, this ability enables the use of the ER pipeline’s output as reference data for subsequent runs. While treating entity profiles as entity references provides flexibility, it also introduces challenges. The traceability and explainability of ER may be reduced, as subsequent links derived from newly generated entity profiles could obscure provenance, affecting auditability and trust. Additionally, if entity clusters lack coherence, erroneous merges may propagate downstream, exacerbating the bad-triplet problem.
@@ 6 ER Runtime Process
Up until now the focus was on generic use cases and abstract logical components. Procedurally ER may be interpreted as a series of subprocesses and activities which, in abstract, remain consistent across existing approaches.
@@ 6.1 Entity Reference Extraction
Entity resolution systems differ in their representation of entity references, which may take the form of tuples or mappings [21, 51], vectors [44, 33], or vertices within a logical graph [58]. Regardless of the chosen format, structuring data is a fundamental step in the ER process. It ensures that the input is represented in a consistent and processable manner.
@@
Once entity references are assembled, the process continues iteratively until no further input data remains. This allows enough flexibility to design batch, progressive, and online ER systems.
@@ 6.4 Entity Profile Assembly
An interesting feature of entity profiles in certain ER models [3] is that they are in turn entity references. This allows for a neat trick of feeding entity profiles back into the system as entity references. The natural place to store profiles is the entity reference store (Subsection 5.5).
@@ 7 Implementation and Deployment Guidelines
Common idioms and patterns emerge from all reference systems even though every implementation is unique.
7.0.1 Development Considerations
The inspected reference systems exhibit idiomatic programming techniques and favor abstraction over hardcoded logic. This approach improves system extensibility, allowing support for diverse data formats and algorithms which drive the ER process. Separating the user interface from the core ER system is beneficial, especially when the UI evolves independently [58, 51]. Similarly, modularizing ER phases enables independent evolution and scalability of components.
@@ 7.0.1 Development Considerations
While the JVM, Python, and R dominate current ER implementations, the problem itself is not bound to any particular technology, as ER models align well with multiple programming paradigms including functional programming, logic programming and object oriented programming.
@@ 7.1 Deployment Considerations
The scalability needs of a system determine in large part its deployment. ER systems scale proportionally to the volume of processed entity references. This is true both of batch and incremental ER systems. Lightweight deployments are feasible for small datasets, but scaling the deployment often presents its set of challenges and trade-offs.
@@ 8.1 Concluding Remarks on Research Questions
To address the research questions posed in this article, the proposed reference architecture for ER is designed with extensibility, interoperability, and scalability as core principles. The architecture proposes modular designs, where components are loosely coupled, allowing for seamless integration of new data sources and matching strategies. To ensure scalability, distributed computing techniques and efficient indexing mechanisms are suggested, reducing computational overhead and improving system responsiveness.
Key logical and process abstractions were identified in Sections 5 and 6 to support both batch and incremental ER. High-throughput is suggested for batch processing pipelines which ensures the processing of large datasets by leveraging parallel computation and efficient blocking techniques. Incremental ER are suggested to handle real-time updates by maintaining stateful representations of resolved entities and dynamically merging new data with minimal reprocessing. Suggestions for implementing hybrid approaches rely on allowing transitions between batch and incremental modes.
The reference architecture addresses the diversity of matching and clustering strategies by considering both of them black boxes, yet places them within a structured framework. Matching pluggability in particular enables the integration of deterministic, probabilistic, machine-learning-based, and other ER models. A similar statement can be made about clustering. The core black boxes at the center of ER are surrounded by logical components which standardise the input and the output of ER. On one side, the comparison space logical abstraction along with the comparison space generation subprocess are introduced to standardize the reduction of candidate entity references used at the core. On the other hand, entity profile assembly is guided by general principles such as the adherence to an ER model to determine their data representation. Additionally, entity profile assembly is comprised from standard activities which do not vary in the abstract across ER systems.
ER configuration and evaluation use cases are further examples of elements left completely open for interpretation. The architecture suggests active learning and human-in-the-loop mechanisms for refining matching rules based on user feedback. Although not explicitly targeted, suggestions are presented for ER evaluation, such as the necessity to evaluate clustering performance in ER.
To increase the interoperability of the implementing system architecture, the reference architecture suggests design patterns useful for standardizing data exchange with users and external systems. Design patterns are also suggested to create more extensible system architectures with respect to the representation of entity profiles or the support of multiple execution engines. But probably the most important design pattern contributed by the reference architecture is the most obvious one: the ER runtime is a pipeline.
Overall, the article presents a structured and adaptable approach to ER, within certain bounds. The presented abstractions and general concepts, as well as the suggested design principles and patterns can constitute a solid foundation for future research and design.
8.2 Feedback Process
This reference architecture is grounded in the analysis of existing open-source ER systems. The sheer amount of existing information makes this initial version of the reference architecture an imperfect work, at best. Nevertheless, by synthesizing best practices observed in systems such as [51, 44, 34, 58], it hopes to provide a structured yet adaptable foundation for designing ER solutions.
The reference architecture should validated primarily through the alignment with the reference implementations, demonstrating that the architecture is not only theoretically sound but also practically applicable. Future refinements should be guided by further empirical evaluation and contributions from the ER research and practitioner communities.
Architects and researchers interested in advancing this reference architecture are encouraged to participate in organizing a framework for collaboration with the goal of evolving it. At a minimum level, the collaboration framework should incorporate feedback from ER implementations in production, real-world constraints, and emerging trends. The goal should be to maintain a living, adaptable design that continues to inform and support the design of bespoke ER system architectures.
@@ 8.3 Final Words
ER is an area of research with a vast and diverse literature which may cause a paradox of choice for newcomers to the field. The present paper introduces a reference architecture for ER focused on extensibility, interoperability and scalability. The reference architecture delineates the bounds where generalization and abstraction yield diminishing returns, particularly for ER configuration and evaluation, but also for matching and clustering to a lesser extent. Yet, to our knowledge, this is the first attempt at writing a reference architecture for ER and the first attempt to unify so many ER flavours using the same abstractions.
The paper was written in hopes of triggering an effort to standardize ER design across all flavours of ER. Both subsequent iterations of this reference architecture, based on feedback and collaboration, and the establishment of an organization to standardize the ER problem domain would be welcome.

### v2

@@ 5.6 Clustering Engine
Clustering plays a crucial role in entity resolution (ER) by grouping linked entity references that correspond to the same real-world object. It follows the principle that things which co-occur or exhibit strong relational ties likely refer to the same underlying entity. This mechanism is complementary to matching which uses the general concept of similarity to provide the evidence of solid relational ties. Both mechanisms contribute to ER in an essential way, though only at least one of them is strictly necessary.
@@
Widely regarded as an essential step, not all reference systems implement it transparently. Some ER systems, such as Duke [21] obfuscate clustering by making it part of intermediate storage operations which use its system of MatcherListener. System developers are informed about clustering possibilities only once the LinkDatabaseMatchListener and LinkDatabase concepts are perused. This is the only example of obfuscating the pipeline structure of the ER process among the reference ER systems. On the other hand, some entity matching systems such as Ditto [35] or DeepMatcher [46] focus solely on identifying matching entity references, partially or completely omitting transitive resolution. In some cases, this approach can be beneficial, as it offloads the bad-triplet problem [1] to downstream systems.
@@
At the other end of the spectrum, entity alignment and record linkage systems designed around graph models inherently support clustering. Some ER solutions go further by eliminating the matching phase entirely. For example, clustering subsumes matching by employing functional similarity measures and a carefully chosen merge function [3], generic distance metrics [71], or probabilistic models which form clusters directly from the comparison space using similarity functions to model colocation (D-blink [40]).
@@
When applied, clustering typically operates on data supplied by the reference store. In cases where matching is omitted, the entity reference store effectively defines the relevant comparison space. Systems such as the JedAI [53] and FAMER [60] already support various clustering algorithms [50, 61], offering flexible options for different ER workflows.
@@
Given its varying adoption across reference systems and the limited research on cluster coherence [70], system architectures may design a clustering engine if their specific problem domain and ER model of choice require one. In object-oriented designs, the Strategy design pattern [19] may be used to encapsulate the various algorithms sitting at the core of the clustering engine.
New system architectures should carefully consider clustering engines that work with incremental ER. Only some clustering algorithms (such as agglomerative clustering) are generally incremental, i.e can be easily carried out using the batch linkage method [24].
Finally, the evaluation of clustering performance is not as common among the reference ER systems as the fundamental importance of the clustering subsystem to ER would imply. D-blink [40] is an example system which implements this feature. ER system architectures which incorporate clustering engines should also provide the means for evaluating clustering performance along with matching performance. Future efforts in this area would improve ER system transparency and reliability.
@@ 5.7 Presentation Layer
The logical component found across all ER systems is the entity profile presentation layer which has the responsibility of making entity profiles available to ER consumers. While entity profile formatting or display vary according to the needs of each implementation, all ER systems define a logical component which assembles entity profiles. The data representation of entity profiles varies along with the chosen approach to ER. The act of designing ER system architectures necessarily involves choosing an ER model which best suits the problem domain. The entity profile assembler logic will necessarily be dictated by the selected ER model.
@@
This poses a problem for the interoperability of ER systems because of the myriad ways to represent entity profiles used in practice. Ditto represents entity profiles as pairs of matching entity references (Ditto [35]). Entity profiles take the form of algebraic set partitions when using Duke’s equivalence class linker [21], JedAI’s ExactClustering component [53] or D-blink’s default output [40]. FAMER represents entity profiles as subgraphs [60]. Similarly, entity alignment [65] systems also represent entity profiles as graphs. The main distinction is that in this case entity profiles are created guided by the knowledge graph’s ontology rather than by relationships between entity references.
SERF [3] introduces a black-box merge function that constructs more dominant entity references incrementally using the matching engine’s output. The R-Swoosh algorithm which defines this approach is exemplified in practice by the Oyster [67] ER system. The final entity references constructed in this way represent the entity profiles, their data structure being determined flexibly by the black box merge function. Merge dominance logically enables the use of the ER pipeline’s output as entity references in subsequent runs. Although treating entity profiles as entity references provides flexibility, it also introduces challenges. The traceability and explainability of ER may be reduced, as subsequent links derived from newly generated entity profiles could obscure provenance, affecting auditability and trust. Additionally, if entity clusters lack coherence, erroneous merges may propagate downstream, exacerbating the bad-triplet problem.
@@
Viewing incremental ER as a query-driven system can hint at possible ways to solve the interoperability problem mentioned earlier in this section. A component oriented design coupled with a web service architecture might provide the standardization required to make ER systems more interoperable.
@@ 6 ER Runtime Process
Up until this point, the paper focused on generic use cases and abstract logical components. Procedurally ER may be interpreted as a series of subprocesses and activities which, in abstract, remain consistent across existing approaches.
@@ 6.1 Entity Reference Extraction
Entity resolution systems differ in their representation of entity references, which may take the form of tuples [46], mappings [21, 36], sets of named attributes [53], vectors [35], or vertices within a logical graph [60]. Regardless of the chosen format, structuring data is a fundamental step in the ER process. It ensures that the input is represented in a consistent and processable manner.
@@
Once entity references are assembled, the process continues iteratively until no further input data remains. This allows enough flexibility to design batch, incremental, and online ER systems.
@@ 6.4 Entity Profile Assembly
Certain ER models (like SERF [3]) allow entity profiles to be viewed as entity references. This allows feeding entity profiles back into the system by storing entity profiles is the entity reference store described in Subsection 5.5.
@@ 7 Implementation and Deployment Guidelines
7.1 Development Considerations
Even though every implementation is unique, the inspected reference systems exhibit idiomatic programming techniques and favor abstraction over hardcoded logic. This approach improves system maintainability, and indirectly extensibility, allowing support for diverse data formats and algorithms which drive the ER process. Separating the user interface from the core ER system is beneficial, especially when the UI evolves independently. FAMER [60] and JedAI [53] exemplify this approach in practice. Similarly, modularizing ER phases enables independent evolution and scalability of components.
Systems that implement a clear and structured pipeline (e.g., FAMER, JedAI, Splink, dblink) are more maintainable than those lacking an idiomatic runtime configuration (e.g., Ditto, DeepMatcher) or those which apply patterns in unconventional ways. For instance, Duke employs the Observer pattern [19], typically used for event-driven programming, to manage workflow execution. We advocate extending the Law of Least Astonishment [26] to source code. In other words, code should be structured in a way that minimizes surprises for reviewers. While this guideline was formulated in a less formal context than literate programming [29], it remains a surprisingly effective principle for improving code maintainability and readability, thereby increasing extensibility as well.
@@ 7.1 Development Considerations
While the JVM, Python, and R dominate current ER implementations, the problem itself is not bound to any particular technology. All ER models can be materialized easily using widespread programming paradigms such as modularization, object orientation or componentization. Our selection of reference systems clearly exemplifies this possibility.
@@ 7.2 Deployment Considerations
The scalability needs of a system determine in large part its deployment. ER systems scale proportionally to the volume of processed entity references. This is true both of batch and incremental ER systems. Lightweight deployments are feasible for small datasets or system development, but scaling the deployment often presents its set of challenges and trade-offs.
@@ 8.1 Concluding Remarks on Research Questions
In the preceding sections, we saw that the main use cases of ER systems broadly fall into three categories: configuration, runtime and evaluation use cases. Configuration use cases were referred to as the ‘design time’ interface of the system and highlighted user roles and potential actions that occur when the ER system is prepared for operation. In contrast, the runtime use cases highlight the activities surrounding ER systems which occur while they resolve entity references into entity profiles. Finally, evaluation use cases highlight activities which are performed either at design time or at runtime to assess the quality and performance of ER.
Then we proceeded by outlining logical (Section 5) and procedural (Section 6) abstractions common to all ER systems, regardless of their focus—record linkage, deduplication, entity alignment, NERD, or other variations. These abstractions apply to both batch and incremental ER. Recommendations were provided, and potential pitfalls addressed to ensure that the logical and procedural components of ER system designs remain extensible, interoperable, and scalable. Key design principles, such as loose coupling and high cohesion, were emphasized alongside established patterns (like the pipeline model) to enhance code maintainability and readability. Distributed computing frameworks, including Apache Spark and Flink, were advocated as preferable alternatives to custom-built runtimes. From the outset, component orientation was recommended, particularly given the black-box nature of match and merge operations.
This brings us to the core question: what key insights can we offer for designing entity resolution systems? The answer to this question determines the practical effectiveness of the reference architecture.
A fundamental requirement for ER systems is end-to-end entity resolution. While we have argued that either matching or clustering is sufficient for ER, systems should strive to incorporate multiple methods for determining which entity references contribute to an entity profile. Our firm recommendation is that ER pipeline components be designed to create value beyond their individual contributions. This is achievable only through careful use of abstractions such as the ER pipeline and the reference store. Further, batch and incremental ER both provide value, but truly remarkable ER systems implement both paradigms. Incremental ER in particular lends itself extremely well to modern architectural approaches such as service-oriented architecture or event-driven architecture.
A natural consequence of end-to-end ER design is the need for appropriate evaluation. Assessing clustering performance solely through statistical metrics (e.g., precision, recall, F-measure) leads to information loss. Conversely, evaluating pairwise clustering performance when no matching has occurred introduces extraneous data with no meaningful interpretation. Regardless of the specific combination of matching, clustering, or other pipeline components, ER evaluation must be implemented and finely tuned to the runtime environment.
Another critical consideration is the flexibility of the reference architecture. Configuration, matching, and clustering remain largely at the discretion of system designers. While this stance may be viewed as a limitation of the reference architecture, avoiding the prescription of rigid implementations for these areas is wholly intentional. Configuration choices are often tied to the underlying technology, while matching and clustering depend on the selected ER model. Furthermore, future methods for selecting entity references beyond matching and clustering may emerge, and the reference architecture strongly supports extensibility in this regard. Matching flexibility allows for deterministic, probabilistic, and machine-learning-based ER models. Clustering approaches may include algebraic, probabilistic, or graph-based methods. The pipeline design enables seamless substitution of rule-based engines with neural network models.
Beyond extensibility, large-scale ER requires careful attention to scalability and performance. Comparison space generation and reference storage are fundamental abstractions which must be present in all ER systems. A component-oriented approach is strongly recommended, as it facilitates scalability by providing standardized interfaces between component-specific logic and the overarching pipeline.
ER systems must efficiently process diverse input data while maintaining strict output consistency. To support downstream integration, the reference architecture promotes standardization: input flexibility is encouraged, while output representation is constrained. Logical components surround the core ER processes, ensuring uniformity in input reduction via comparison space generation and in entity profile assembly based on ER models. Despite differences in implementation, core entity profile assembly activities remain consistent across ER systems.
These principles lead to a reference architecture that prioritizes component-oriented design, emphasizing extensibility in configuration, core runtime components, and evaluation. Interoperability is achieved by allowing input flexibility through entity reference extraction while enforcing output standardization through entity profile assembly. Scalability and performance are upheld through essential abstractions, including comparison space generation and the reference store.
Ultimately, this article presents a structured yet adaptable approach to ER within well-defined boundaries. The proposed abstractions, principles, and patterns combine themselves into a reference architecture which can be a structured foundation for future research and practical implementations.
8.2 Feedback Process and Validation
The reference architecture proposed in this paper was developed through a rigorous study and synthesis of existing systems. By abstracting their source code into generic concepts and processes, it ensures empirical soundness. Efforts were made to maintain a clear correspondence between these abstractions and their concrete implementations, as outlined throughout the paper. As such, the architecture provides a structured yet flexible foundation for designing ER solutions. However, given the vast scope of existing knowledge, this initial version remains an imperfect work, at best.
Rather than developing another ER system (which would only contribute to the paradox of choice aluded to in the introduction), a more effective approach is to iteratively refine the architecture based on feedback from practitioners. This transforms it into a dynamic framework, continuously improved using empirical evaluation and contributions from the ER community.
To support this evolution, researchers and architects are encouraged to reach out to the authors to establish a collaborative framework. At a minimum, this approach would allow integrating insights from real-world implementations, industry constraints, and emerging trends. The ultimate goal is to sustain an adaptive design that informs and supports the development of ER systems over time.
@@ 8.3 Final Words
ER is an area of research with a vast and diverse literature which may cause a paradox of choice for newcomers to the field. The present paper introduces a reference architecture for ER focused on extensibility, interoperability and scalability. The reference architecture delineates the bounds where generalization and abstraction yield diminishing returns, particularly for ER configuration and evaluation, but also for the ER runtime to a lesser extent. Yet, to our knowledge, this is the first attempt at writing a reference architecture for ER and the first attempt to unify so many ER flavours using the same abstractions.
The paper was written in hopes of triggering an effort to standardize ER design across all flavours of ER. Subsequent iterations of this reference architecture (based on feedback and collaboration), and the standardization of the broader ER problem domain would both be welcome.